A manufacturer produces lots of a canned food product. Let p denote the proportion of the lots that do not meet the product quality specifications. An

n = 31, c = 0 acceptance sampling plan will be used.

(a) Compute points on the operating characteristic curve when p = 0.01, 0.03, 0.10, and 0.20. (Round your answers to four decimal places.)

c

p = 0.01

p = 0.03

p = 0.10

p = 0.20

0

(b) Plot the operating characteristic curve.

(c) What is the probability that the acceptance sampling plan will reject a lot containing 0.03 defective? (Round your answer to four decimal places.)
Answer #1

a)

for 0.01 ; probability of accepting the lot =(1-0.01)^31 =0.7323

from above:

p P(accepting the lot)
0.01 0.7323
0.03 0.3890
0.1 0.0382
0.2 0.0010

b)

Placcepting the lot) 0.8000 0.7000 0.6000 0.5000 0.4000 0.3000 0.2000 0.1000 0.0000 0 0.05 0.1 0.15 0.2 0.25

c)

P(reject) =1-0.3890 =0.6110